Grade 12th passMechanics

A body of mass 2kg makes an elastic collision with another body at rest and continues to move in the original direction at a speed equal to 1/3 of its original speed. The mass of the second body is

swathi
8 Years agoGrade 12th pass

1 Answer

Bhavasagar
8 Years ago
According to law of conservation of energy,m1u1+m2u2=m1v1+m2v2. Let,v1=v. (2)v+(m2)0=(2)(1/3v)+(m2)(v2). 2(v)=2/3v+(m2)(v2). m2(v2)=4/3v. In elastic collision kinetic energy is also conserved,so. 1/2(2)v^2+1/2(m2)(0)=1/2(2)(1/3v)^2 +1/2(m2)(v2)^2...... 1/2(m2)(v2)^2=8/9(v^2)..... (m2)(v2)^2=16/9v^2.... (m2)(v2)(v2)=16/9v^2..... (v2)=4/3(v)... On substituting v2 in,. 2v+0=2/3v + (m2)(4/3v). (Since v2=4/3v). (m2)=1kg ......